OpenStudy (anonymous):

Which of the following is not an element of the following set: A: {negative even integers} −4 −18 −100 −23

OpenStudy (anonymous):

Well, you're looking for whichever number doesn't divide evenly by 2.

OpenStudy (anonymous):

d?

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

wow.. I cant belive Im getting them right:D right on!

OpenStudy (anonymous):

\[-\frac{4}{2}=-2\] \[-\frac{18}{2}=-9\] \[-\frac{100}{2}=-50\]

OpenStudy (anonymous):

If N = {all real numbers} and subset S = {multiples of 2 greater than zero}, what is S′? {all real numbers excluding multiples of 2 greater than zero} {multiples of 2 less than zero} {even numbers greater than zero} {odd numbers greater than zero}